OpenStudy (anonymous):

Challenge: Find the radius of the base of a cone with volume equal to 144π in3 and height equal to 12 in. A. 6 in B. 3.8 in C. 36 in D. 9 in

OpenStudy (goldphenoix):

To find the volume of a cone, you use the formula: \[\large \large \frac{ 1 }{ 3 } \pi r^2(12) = 144\pi\]\] Now we are given the volume, and the height. So we just plug it in:

OpenStudy (goldphenoix):

\[\large \large \frac{ 1 }{ 3 } \pi r^2(12) = 144\pi\]

OpenStudy (goldphenoix):

First off, we want to get rid of pi in the right side. So 144/ pi give us 144. If we divide pi on the right side, then we must do it to the left side as well. So what does pi divided by pi give us?

OpenStudy (anonymous):

1

OpenStudy (goldphenoix):

Great. So that would cancel out. Now we are left with: \[\large \large \frac{ 1 }{ 3 } r^2h(12) = 144\] What can we simplify in the left side?

OpenStudy (goldphenoix):

My bad about the equation. Take out the h.

OpenStudy (anonymous):

wouldnt it be 4hr^2 or something like that then?

OpenStudy (goldphenoix):

It would be 4 * 12 * r^2

OpenStudy (goldphenoix):

Wait.

OpenStudy (goldphenoix):

Arg. I did something wrong. But I have a feeling that it's A.

OpenStudy (anonymous):

i worked through it as much as i could, and i just couldnt come up with any that nmatched, i don't know what i was doing wrong either. im super unfamiliar with thgis stuff. but it was A, so thanks for the better guess than me.